9. A crane is to lift 2000 kg of material to a height of 150 m in 1 min at a uniform rate. What electric power is required to drive the crane motor if 35 percent of the electric power is converted to mechanical power? (is it reasonable to neglect the kinetic energy in this process?)
Answer #1

given
m = 2000 kg
h = 150 m
t = 1 min = 60 s

wer know,
mechanical power required, P = Workdone/time taken

= m*g*h/t

= 2000*9.8*150/60

= 49000 W

35% of electrical power = mechanical power

0.35*electrical power = mechanical power

electrical power = mechanical power/0.35

= 49000/0.35

= 140000 W <<<<<<<<<------------------Answer